# ] 1NLPR, MAIS, CASIA  2UCAS  3ByteDance

### v1

@@ 4 TreeVGR
Notably, our TreeVGR does not require actually replaying cropped images as previous approaches [60, 76, 50] do, as text-space grounding is already effective. It leads to much more efficient training and inference procedures.
@@ 4.1 Cold-Start Initialization
While end-to-end reinforcement learning (RL) has demonstrated validity by [76] for visual grounded reasoning (VGR) tasks, its practical deployment remains hindered by extreme computational demands. Specifically, DeepEyes-7B [76] requests RL training on 47K samples across 32 episodes, a process requiring 32 H100 (80GB) GPUs operating continuously for 48 hours. Such resource intensity creates barriers to iterative experimentation and broader accessibility.
To address these limitations, we investigate a computationally efficient alternative. Initial attempts revealed significant training inefficiencies when applying direct RL to VGR: models required extensive iterations to autonomously identify task-relevant visual regions before generating answers. This bottleneck motivates our adoption of a cold initialization strategy as illustrated in Figure 7a. Specifically, we introduce a supervised fine-tuning (SFT) phase using a curated dataset comprising multimodal samples: each sample includes an image, a question, reasoning trajectories with corresponding bounding boxes, and a final answer. This structured initialization ensures VGR capabilities are established prior to RL.
Data Construction. We base our supervised fine-tuning (SFT) dataset on VGR-158K [60], which provides pseudo-chain-of-thought annotations paired with bounding boxes for visual reasoning tasks. However, to align with the grounding capabilities of our base model (Qwen2.5-VL series [2]), which outputs absolute coordinates rather than the normalized coordinates (ranging from 0 to 1) used by LLaVA-NeXT [29] in [60], we perform coordinate system conversion. Specifically, for each bounding box, we transform normalized coordinates $[r_{x_{1}},r_{y_{1}},r_{x_{2}},r_{y_{2}}]$ into absolute coordinates via $[x_{1},y_{1},x_{2},y_{2}]=[Wr_{x_{1}},Hr_{y_{1}},Wr_{x_{2}},Hr_{y_{2}}]$, where $H\times W$ is the resolution of the input image. Next, we filter samples to prioritize complex reasoning pathways, retaining only entries with multiple bounding boxes (i.e., more than one box per reasoning trajectory). This yields 35K samples, as multi-box interactions demand stronger spatial-temporal reasoning compared to single-box tasks. Subsequently, we construct a reflective subset of 4.7K samples among them by introducing controlled perturbations: for each sample, we (1) inject a synthetic error by inserting a randomly generated incorrect bounding box into the reasoning sequence, and (2) append the meta-cognitive prompt “Wait, this box seems to be wrong” immediately afterward, resulting in our TreeVGR-SFT-35K. This design explicitly trains the model to detect and correct erroneous visual grounding, which is a critical skill for robust real-world deployment.
Optimization. Initialized from Qwen2.5-VL-7B-Instruct [2], we train TreeVGR-7B-CI (“CI” here stands for Cold Initialization) with 8 GPUs using LLaMA-Factory [74], where the AdamW optimizer [36] with a learning rate of 5e-6 and a global batch size of 256 is utilized. The learning rate is decayed following a cosine schedule [35] with a warmup ratio of 0.1.
@@ 4.2 Reinforcement Learning with Traceable Evidence
where the accuracy reward assesses whether the final answer is correct. We utilize exact-matching for multiple-choice questions, and leverage an online reward model, i.e., Qwen2.5-72B-Instruct [52], to judge whether the prediction is correct given the question and the ground-truth answer. The formatting reward ensures the reasoning process and the final answer must be enclosed between <think> and </think>, and <answer> and </answer>, respectively. The dual IoU reward measures the quality of predicted boxes against ground-truths. Specifically, for $N$ predicted bounding boxes $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$, where $\hat{\bm{b}}_{i}=[\hat{x}_{1}^{i},\hat{y}_{1}^{i},\hat{x}_{2}^{i},\hat{y}_{2}^{i}]$ and $M$ ground-truths $\{\bm{b}_{k}\}_{k=1}^{M}$, where $\bm{b}_{k}=[x_{1}^{k},y_{1}^{k},x_{2}^{k},y_{2}^{k}]$, the dual IoU is an average of a recall term and a precision term.
@@
where the first term $R_{\text{IoU}}^{\text{R}}$ indicates the recall and the second term $R_{\text{IoU}}^{\text{P}}$ means the precision. Specifically, the recall term ensures that each ground-truth bounding box $\bm{b}_{k}$ is matched with at least one prediction.
@@
Similarly, $\text{IoU}\left[\{\bm{b}_{k}\}_{k=1}^{M},\hat{\bm{b}}_{i}\right]=\max_{k}\text{IoU}(\bm{b}_{k},\hat{\bm{b}}_{i})$ indicates the maximum IoU between all ground-truths $\bm{b}_{k}$ and each prediction $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$. Maximizing this term encourages each prediction $\hat{\bm{b}}_{i}$ to be matched with at least one ground-truth. Therefore, simultaneous optimization of both recall and precision eliminates the need for exhaustive enumeration of bounding boxes, thereby contributing to more accurate reasoning pathways.
@@
Gemini-2.5-Flash-0520 [9] 45.9 48.3 53.9 69.6 68.8 75.0 15.3 19.3 56.1 72.4 43.2
GPT-4o-1120 [41] 46.9 51.7 61.5 65.2 43.8 69.1 18.8 38.6 48.8 72.4 43.2
Gemini-2.5-Pro-0605 [10] 54.1 51.7 61.5 56.5 75.0 83.8 20.0 36.8 65.9 86.2 54.6
o3-0416 [43] 54.8 69.0 69.2 65.2 68.8 79.4 22.4 38.6 61.0 86.2 50.0
@@
LLaVA-OneVision-7B [23] 37.3 55.2 53.8 56.5 50.0 32.4 21.2 22.8 41.5 72.4 36.4
LLaVA-OneVision-72B [23] 40.5 62.1 53.8 65.2 62.3 36.8 12.9 28.1 53.7 65.5 47.7
Qwen2.5-VL-7B [2] 37.0 55.2 53.8 56.5 62.5 27.9 20.0 35.1 39.0 44.8 43.2
Qwen2.5-VL-32B [2] 42.5 51.7 53.8 69.6 62.5 54.4 16.5 33.3 46.3 62.1 38.6
Qwen2.5-VL-72B [2] 42.2 65.5 69.2 56.5 56.3 48.5 11.8 33.3 51.2 72.4 38.6
InternVL3-8B [77] 38.8 51.7 69.2 56.5 56.3 33.7 21.2 24.6 39.0 72.4 43.2
InternVL3-38B [77] 42.0 51.7 61.5 52.2 68.8 51.5 12.9 33.3 56.1 65.5 38.6
InternVL3-78B [77] 46.4 62.1 61.5 52.2 68.8 52.9 16.5 33.3 61.0 86.2 45.5
@@
DeepEyes-7B [76] 37.5 30.0 62.1 53.8 65.2 68.8 51.5 11.8 24.6 36.6 51.7 47.7
Pixel-Reasoner-7B [50] 39.0 35.7 58.6 61.5 65.2 50.0 48.5 14.1 31.6 39.0 44.8 40.9
@@
Table 2: Selected results of different models on TreeBench. Evaluations of open-source general models are implemented using VLMEvalKit [12], while evaluations of visual grounded reasoning models are conducted by us. Reasoning pathways of o3 [43] are unavailable, and thus traceable evaluations are not valid. Best performances for open-source models are highlighted in bold. Our TreeVGR-7B achieves comparable performance with InternVL3-78B [77].
Data Construction. As discussed above, TreeVGR incorporates a novel dual IoU reward, which means each sample should contain ground-truth bounding boxes during the RL phase. To this end, we filter hard samples from the original 191K training set of V* [66] using Qwen2.5-VL-7B-Instruct [2], resulting in 30K samples. Additionally, we incorporate the VisDrone dataset [78], which is originally designed for detection and tracking under UAV images, which offers extremely high-resolution, real-world scenes with a large number of small and varied objects and their corresponding bounding box annotations. We reformulate the training set and the validation set into 38K multiple-choice counting problems, and only retain samples with the ground-truth number ranging from 5 to 10, contributing to the final 7K samples. Finally, our TreeVGR-RL-37K consists of 30K open-ended question-answering samples from V* [66] and 7K multiple-choice problems from VisDrone [78].
Optimization. Initialized from TreeVGR-7B-CI, we train our final TreeVGR-7B with 8 GPUs, with another 8 GPUs serving the reward model, i.e., Qwen2.5-72B-Instruct [52], using vLLM [21]. We adopt Group Relative Policy Optimization (GRPO) [47], which has been proved to be effective and efficient for diverse tasks. We have also tried DAPO [70], but we find it unstable compared with GRPO. Therefore, we simply utilize the original GRPO [47]. We implement using EasyR1 [75], which is a clean fork of veRL [49]. We train our TreeVGR-7B with 5 epochs on TreeVGR-RL-37K, which is significantly less than DeepEyes-7B [76] (which is trained on 47K samples with 32 epochs).

### v2

@@ 4 TreeVGR
Notably, our TreeVGR does not require actually replaying cropped images as previous approaches (Wang et al., 2025e; Zheng et al., 2025b; Su et al., 2025) do, as text-space grounding is already effective. It leads to much more efficient training and inference procedures.
@@ 4.1 Cold-Start Initialization
While end-to-end reinforcement learning (RL) has demonstrated validity by (Zheng et al., 2025b) for visual grounded reasoning (VGR) tasks, its practical deployment remains hindered by extreme computational demands. Specifically, DeepEyes-7B (Zheng et al., 2025b) requests RL training on 47K samples across 32 episodes, a process requiring 32 H100 (80GB) GPUs operating continuously for 48 hours. Such resource intensity creates barriers to broader accessibility.
To address these limitations, we investigate a computationally efficient alternative. Initial attempts revealed significant training inefficiencies when applying direct RL to VGR: models required extensive iterations to autonomously identify task-relevant visual regions before generating answers. This bottleneck motivates our adoption of a cold initialization strategy as illustrated in Figure 3a. Specifically, we introduce a supervised fine-tuning (SFT) phase using a curated dataset comprising multimodal samples: each sample includes an image, a question, reasoning trajectories with corresponding bounding boxes, and a final answer. This structured initialization ensures VGR capabilities are established prior to RL. Details of data construction and optimization can be found in Appendix E.1.
@@ 4.2 Reinforcement Learning with Traceable Evidence
where the accuracy reward assesses whether the final answer is correct. We utilize exact-matching for multiple-choice questions, and leverage an online reward model, i.e., Qwen2.5-72B-Instruct (Team, 2024), to judge whether the prediction is correct given the question and the ground-truth answer. The formatting reward ensures the reasoning process and the final answer must be enclosed between <think> and </think>, and <answer> and </answer>, respectively. The dual IoU reward measures the quality of predicted boxes against ground-truths. Specifically, for $N$ predicted bounding boxes $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$, where $\hat{\bm{b}}_{i}=[\hat{x}_{1}^{i},\hat{y}_{1}^{i},\hat{x}_{2}^{i},\hat{y}_{2}^{i}]$ and $M$ ground-truths $\{\bm{b}_{k}\}_{k=1}^{M}$, where $\bm{b}_{k}=[x_{1}^{k},y_{1}^{k},x_{2}^{k},y_{2}^{k}]$, the dual IoU is an average of a recall term and a precision term.
@@
where the $R_{\text{IoU}}^{\text{R}}$ indicates the recall and $R_{\text{IoU}}^{\text{P}}$ means the precision. Specifically, the recall term ensures that each ground-truth bounding box $\bm{b}_{k}$ is matched with at least one prediction.
@@
Similarly, $\text{IoU}\left[\{\bm{b}_{k}\}_{k=1}^{M},\hat{\bm{b}}_{i}\right]=\max_{k}\text{IoU}(\bm{b}_{k},\hat{\bm{b}}_{i})$ indicates the maximum IoU between all ground-truths $\bm{b}_{k}$ and each prediction $\{\hat{\bm{b}}_{i}\}_{i=1}^{N}$. Maximizing this term encourages each prediction $\hat{\bm{b}}_{i}$ to be matched with at least one ground-truth. Therefore, simultaneous optimization of both recall and precision eliminates the need for exhaustive enumeration of bounding boxes, thereby contributing to more accurate reasoning pathways. Details of data and optimization can be found in Appendix E.2.
Table 2: Selected results of different models on TreeBench. Evaluations of open-source general models are implemented using VLMEvalKit (Duan et al., 2024), while evaluations of visual grounded reasoning models are conducted by us. Reasoning pathways of o3 (OpenAI, 2025) are unavailable, and thus traceable evaluations are not valid. Best performances for open-source models are highlighted in bold. Our TreeVGR-7B achieves comparable performance with InternVL3-78B (Zhu et al., 2025).
@@
Gemini-2.5-Flash-0520 45.9 48.3 53.9 69.6 68.8 75.0 15.3 19.3 56.1 72.4 43.2
GPT-4o-1120 46.9 51.7 61.5 65.2 43.8 69.1 18.8 38.6 48.8 72.4 43.2
Gemini-2.5-Pro-0605 54.1 51.7 61.5 56.5 75.0 83.8 20.0 36.8 65.9 86.2 54.6
o3-0416 54.8 69.0 69.2 65.2 68.8 79.4 22.4 38.6 61.0 86.2 50.0
@@
LLaVA-OneVision-7B 37.3 55.2 53.8 56.5 50.0 32.4 21.2 22.8 41.5 72.4 36.4
LLaVA-OneVision-72B 40.5 62.1 53.8 65.2 62.3 36.8 12.9 28.1 53.7 65.5 47.7
Qwen2.5-VL-7B 37.0 55.2 53.8 56.5 62.5 27.9 20.0 35.1 39.0 44.8 43.2
Qwen2.5-VL-72B 42.2 65.5 69.2 56.5 56.3 48.5 11.8 33.3 51.2 72.4 38.6
InternVL3-8B 38.8 51.7 69.2 56.5 56.3 33.7 21.2 24.6 39.0 72.4 43.2
InternVL3-78B 46.4 62.1 61.5 52.2 68.8 52.9 16.5 33.3 61.0 86.2 45.5
@@
DeepEyes-7B 37.5 30.0 62.1 53.8 65.2 68.8 51.5 11.8 24.6 36.6 51.7 47.7
Pixel-Reasoner-7B 39.0 35.7 58.6 61.5 65.2 50.0 48.5 14.1 31.6 39.0 44.8 40.9
@@ 5 Experiments
Baselines. We include four state-of-the-art private models, GPT-4o-1120 (OpenAI, 2024a) and o3-0416 (OpenAI, 2025) from OpenAI, and Gemini-2.5-Flash-0520 (DeepMind, 2025a) and Gemini-2.5-Pro-0605 (DeepMind, 2025b) from Google. Additionally, representative open-source general models are incorporated, including LLaVA-OneVision series (Li et al., 2024), Qwen2.5-VL series (Bai et al., 2025a), and InternVL3 series (Zhu et al., 2025). Furthermore, two very recent visual grounded reasoning models are also included, i.e., DeepEyes (Zheng et al., 2025b) and Pixel-Reasoner (Su et al., 2025), as both of them follow a “grounding then answering” pipeline, with the capability of “thinking with images”. Evaluations are mainly conducted on TreeBench, V* Bench (Wu and Xie, 2024), HR-Bench (Wang et al., 2025f), and MME-RealWorld-Lite (Zhang et al., 2024a).
Results on TreeBench. Table 2 presents per per-category performance of different models. Overall, OpenAI’s o3-0416 (OpenAI, 2025), the state-of-the-art visual grounded reasoning model, demonstrates the strongest perception abilities, as expected. Larger models usually contribute to better performance. Notably, our TreeVGR-7B even achieves comparable performance with InternVL3-78B (Zhu et al., 2025), demonstrating the effectiveness of the visual grounded reasoning pipeline. Moreover, compared with visual grounded reasoning models, our TreeVGR not only achieves a higher overall performance, but also obtains a larger mIoU, indicating its effectiveness in precisely localizing target objects. More in-depth analysis on TreeBench can be found in Appendix D.
Table 3: Comparison with state-of-the-art alternatives on V* Bench (Wu and Xie, 2024) and HRBench (Wang et al., 2025f). All results are self-collected. Best performances of visual grounded reasoning models are highlighted in bold.
V* Bench HR-Bench-4K HR-Bench-8K
Overall Attr. Spatial Overall Single Cross Overall Single Cross
Private Models
GPT-4o-1120 66.0
o3-0416 95.7
Open-source General Models
LLaVA-OneVision-7B 70.7 73.0 60.5 64.3 74.8 53.8 59.8 65.3 54.3
LLaVA-OneVision-72B 73.8 80.9 63.2 66.3 76.5 56.0 60.9 68.8 53.0
InternVL3-8B 72.3 73.0 71.1 70.8 79.3 62.3 62.0 64.3 59.8
InternVL3-78B 76.4 75.7 77.6 75.5 84.5 66.5 67.3 71.8 62.8
Qwen2.5-VL-7B 74.3 77.4 69.7 72.1 88.8 55.5 68.8 83.5 54.0
Qwen2.5-VL-72B 84.8 90.8 80.9 79.4 88.8 70.0 76.3 84.3 68.3
Open-source Visual Grounded Reasoning Models
Pixel-Reasoner-7B 80.6 83.5 76.3 72.9 86.0 60.3 66.9 80.0 54.3
DeepEyes-7B 90.0 92.1 86.8 75.1 91.3 59.0 72.6 86.8 58.5
TreeVGR-7B 91.1 94.0 87.0 77.1 90.3 64.0 73.1 86.5 59.8
Table 4: Comparison with state-of-the-art alternatives on MME-RealWorld-Lite (Zhang et al., 2024a). All results are self-collected. The best performance is highlighted in bold.
Perception Reasoning
Overall OCR RS DT MO AD OCR DT MO AD
General Models
Qwen2.5-VL-7B 42.3 87.6 32.7 83.0 27.3 30.0 72.0 62.0 28.7 23.0
Qwen2.5-VL-72B 43.7 90.8 34.0 87.0 27.9 30.6 74.0 61.0 26.7 25.5
LLaVA-OneVision-7B 43.7 80.0 40.0 56.0 31.7 39.4 65.0 33.0 38.0 32.0
LLaVA-OneVision-72B 48.7 79.2 50.7 67.0 37.9 40.0 76.0 41.0 38.7 39.3
InternVL3-8B 47.9 83.6 49.3 75.0 34.5 36.9 70.0 44.0 40.0 37.0
InternVL3-78B 52.3 87.6 54.7 77.0 42.6 36.6 76.0 56.0 46.0 40.3
Visual Grounded Reasoning Models
Pixel-Reasoner-7B 49.7 89.6 52.0 86.0 38.9 30.9 71.0 72.0 46.0 32.5
DeepEyes-7B 53.2 90.0 52.7 89.0 43.3 33.4 76.0 69.0 44.0 35.0
TreeVGR-7B 54.9 87.6 50.7 83.0 47.0 43.4 74.0 66.0 51.3 39.0
Results on High-Resolution Benchmarks. In Table 3, TreeVGR achieves open-source state-of-the-art on V* Bench (Wu and Xie, 2024). On HR-Bench (Wang et al., 2025f) and MME-RealWorld-Lite (Zhang et al., 2024a) illustrated in Table 3 and Table 4, respectively, our TreeVGR brings significant improvements over our base model, Qwen2.5-VL-7B (Bai et al., 2025a). Results on other general benchmarks can be found in Appendix F.1.
Ablation Studies. The core contribution of TreeVGR is the traceable training pipeline, where $R_{\text{IoU}}$ is incorporated in conventional RL training. The effectiveness of this design is ablated in Appendix F.2.
6 Conclusion
TreeBench addresses critical gaps in existing benchmarks by focusing on three principles: focused visual perception (identifying subtle targets in cluttered scenes), traceable evidence (quantifiable reasoning chains via bounding box annotations), and vision-centric second-order reasoning. Constructed through expert-driven annotation and multi-stage quality control, TreeBench features 405 high-difficulty visual question-answer pairs with precise bounding boxes, emphasizing small objects in real-world scenarios. It reveals the limitations of state-of-the-art models, e.g., OpenAI-o3 (OpenAI, 2025) scores 54.8%, while setting a new standard for assessing nuanced visual grounding, multi-step reasoning transparency, and cross-modal interaction.
TreeVGR advances VGR training through reinforcement learning guided by dual IoU rewards, which explicitly supervise bounding box generation to ensure both precision and recall. This approach enables explainable reasoning pathways and achieves significant improvements across benchmarks.
Acknowledgements
This work was supported by the Beijing Natural Science Foundation (No. L257015) and the National Natural Science Foundation of China (No. 62320106010).
Ethics Statement
Our research is grounded in ethical practices, with particular attention paid to the responsible use of data. All datasets employed in this study are publicly available and well-established within the computer vision community. Specifically, our benchmarking was conducted on SA-1B (Kirillov et al., 2023). Our use of this data is in accordance with their provided licenses and intended academic purpose.
Reproducibility Statement
We are committed to ensuring the reproducibility of the research presented in this paper. To this end, comprehensive implementation details for our models and experiments are provided in Appendix E, including the training procedures and all hyperparameters used. Furthermore, upon acceptance of this paper, all source code, datasets, and trained model checkpoints will be made publicly available.
Appendix
Appendix A Overview
Here, we provide a table of contents:
  • First, in Appendix B, we provide the annotation pipeline in detail, which includes three rounds of quality control.
  • In Appendix C, we introduce statistics of our TreeBench.
  • In Appendix D, we perform in-depth analysis on our TreeBench.
  • In Appendix E, we provide implementation details of our two-stage training pipeline, including cold-start initialization and reinforcement learning with traceable evidence.
  • In Appendix F, we provide more experiments of our TreeVGR, including results on general multimodal benchmarks and ablation studies.
  • In Appendix G, we discuss our limitations in detail.
  • Finally, in Appendix H, we provide qualitative examples and failure cases of our TreeVGR.
Appendix B Annotation Pipeline
TreeBench was constructed through a systematic pipeline combining automated sampling, LMM-assisted generation, and three rounds of human validation. The annotation team contains eight human experts in LMMs, including six Ph.D candidates and two senior research scientists.
1. Image Selection. A total of 1K images are initially sampled from the SA-1B (Kirillov et al., 2023), with deliberate prioritization of images containing high-density objects (e.g., scenes with overlapping or clustered items), as it offers high-resolution, real-world scenes with a large number of small and varied objects, making it particularly suitable for evaluating visual grounded reasoning. To ensure balanced representation across categories, 100 images are initially allocated per category.
2. First Round Quality Control. The annotation team manually evaluates the relevance and quality of each image for its assigned category. This step is critical for addressing category-specific requirements, e.g., the “Ordering” category necessitates images with visually similar or repetitive objects for practical reasoning tasks. Following this review, 647 images meet the criteria.
3. Automated Question Generation. Question-option-answer trios are then generated using two advanced LMMs, i.e., OpenAI-o3 (OpenAI, 2025) and Gemini-2.5-Pro (DeepMind, 2025b), each tasked with producing three diverse, high-quality questions per image. Prompts are designed to emphasize task-specific complexity and visual-semantic alignment.
4. Second Round Quality Control. Human experts then manually review all six model-generated questions per image. For each image, annotators selected the most semantically coherent and task-relevant question from the pool of six, prioritizing: (1) alignment with the target subtask, (2) avoidance of trivial or ambiguous object referring, and (3) clarity and unambiguous answerability. If none of the six questions met these criteria, annotators manually constructed a new question. This step ensures that only high-quality, human-vetted questions advance to the next stage.
5. Difficulty Filtering. Questions deemed insufficiently challenging are removed through model-based consensus screening. Specifically, any question answered correctly by all four state-of-the-art vision-language models (Qwen2.5-VL-72B (Bai et al., 2025a), InternVL3-78B (Zhu et al., 2025), GPT-4o (OpenAI, 2024a), Gemini-2.5-Flash (DeepMind, 2025a)) was excluded to ensure the benchmark retained meaningful difficulty.
6. Third Round Quality Control. The final cross-verification phase engages independent human annotators to cross-validate the accuracy and relevance of each question-option-answer pair. The final dataset comprised 405 rigorously validated questions.
Appendix C Benchmark Statistics
Distribution of Target Instance Area. We compute the relative area for each target instance using its bounding box, i.e., $\text{area}=\frac{1}{HW}(y_{2}-y_{1})(x_{2}-x_{1})$, where $H$ and $W$ are the input resolution. Figure 7 is the histogram of the mean area for each question. It illustrates that the majority of target instances in TreeBench are extremely small, with a sharp peak near 0.0 and a long tail extending to larger areas (up to 0.7). The mean area across all questions is 0.0305, confirming that targets are predominantly tiny. Most questions (highest frequency bin) involve target instances with areas clustered around 0.0 to 0.05, while only a small fraction require identifying larger objects. This distribution highlights the importance of addressing challenging scenarios where small-scale object detection and reasoning are crucial, potentially compromising model performance.